The area enclosed between the curves y2 = x and y = | x | is ?

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Curves intersect at only two point whose coordinates are:
(0, 0), (1,1)
Area ‘A’ b/w curves:
A = \int_{0}^{1}(\sqrt{x} - x)dx
A = (\frac{2x^{3/2}}{3}-\frac{x^{2}}{2})_{0}^{1} = \frac{1}{6}
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty